# Data Matters:

### v1

@@ V-A3 Machine learning model
Our model has a single LSTM layer with 16 cells. In the training phase, we use the root mean square propagation  and mean squared error (MSE) loss for optimizing the models parameters. We set the length of the historical traffic series to six samples (30 min) due to the speed with which changes in the traffic volume occur in 5-minute intervals. The prediction horizon is the next 5-minute interval. The 24 weeks of data are split 12:6:6 into training, validation, and testing. Training data is shuffled.
V-B How much can road traffic data help?
V-B1 Metrics
We measure the forecasting performance with the mean absolute error (MAE), mean absolute percentage error (MAPE), MSE and root mean squared error (RMSE).
TABLE II: Prediction Performance on Calls Datasets and Flow, Speed and Calls Dataset
@@ V-B1 Metrics
TABLE III: Prediction Performance on New, Handover, and Total Calls Dataset and the Same Dataset but with Flow and Speed Time-Series
@@
TABLE IV: Prediction Performance with Gaussian Noise Added to the Flow
@@
V-B2 Overall performance
@@
The results from the four different data sets—1)  calls (denoting total number of calls, the sum of new and handover calls) only; 2) flow, speed and calls; 3) new, handover, calls and 4) flow, speed, new, handover, calls—when the call duration is modeled by a mixture of two log-normal distributions are summarized in Table II and Table III. For each error metric we show the minimum, median and maximum of the 5 simulation runs conducted for each data set and BS. We make the observations as follows:
@@
  • Road traffic data consistently improve the mobile cellular traffic prediction performance. All BSs observe improvements in all their forecasting performance measures when learning from both network and road data.
@@
    The prediction error reduction is between $8.4\%$ and $33.9\%$ (MAE), $13.3\%$ and $56.5\%$ (MSE), $12.5\%$ and $38.4\%$ (MAPE), and $6.9\%$ and $34\%$ (RMSE) when in addition to (total number of) calls, also flow and speed are employed, and when comparing the medians of all BSs. Similar percentage decrease is recorded for the minimum and maximum errors. Using more refined network KPIs—those that contain the number of new and handover calls, and their total—and road data, reduces the prediction errors too: between $7.5\%$ and $21.8\%$ (MAE), $11\%$ and $35.4\%$ (MSE), $10.9\%$ and $25.7\%$ (MAPE), and $5.7\%$ and $19.7\%$ (RMSE) for the medians and likewise for the smallest and largest errors.
@@
  • The road traffic metrics can capture the processes underlying the mobile cellular traffic generation. We employ flow and speed as a means to characterize the underlying population dynamics and through them the data generation process. The flow captures the number of potential call generation sources and the speed can serve as an indicator for road traffic congestion and hence, for increased traffic load on the BS. The number of handovers (or equivalently, the handover rate) is an indicator of mobility too yet our results are not conclusive about its impact. Contrasting the performance of calls to that of new, handover, and calls, we observe an error improvement in some of the error measures: it is in the order of no more than $3\%$ for the majority of the BSs.
@@
    Finally, employing flow, speed and calls (3 variables) leads to a large reduction in the prediction errors even when contrasted with employing new, handover and calls (3 variables): $8.6\%$ and $22.8\%$ (MAE), $14.2\%$ and $36.3\%$ (MSE), $11.5\%$ and $25\%$ (MAPE), and $7\%$ and $20.2\%$ (RMSE) for the medians.
  • Choice of data (split) and hyper-parameter tuning.
@@
    We highlight that the hyper-parameter tuning is an indispensable part of every machine learning process. In the tables above we show the results from using 24 weeks of data and 75:25 ratio of train and validation to test data. This amount of data seems to be beneficial mainly to the BSs with long range (320280 and 320287) but less to the reaming BSs, which evidence improved performance when trained over shorter periods. We also observe cases when the performance of some of the BSs is decreased when using road data. This is the case for the MSE and RMSE of BS 320287 when the data set consists of weeks 15 to 21 and the data is chronologically split  into 3:2:2 for training, validation and testing [^2] See https://github.com/nvassileva/DataMatters for detailed results.. We identify week 15 as outlying from the other training weeks and consequently, as the main explanation of the results. For the same period but with 4:1:3 partition of the data, the results are consistent with what we report above—decreased prediction errors across all performance measures and BSs. This 4:1:2 data split also turns into smaller forecasting errors for the purely network data sets. In other words, choice of data that captures the underlying patterns and an appropriate hyper-parameter tuning are always required. In our on-going work we are studying the effect of the data split on the overall forecasting performance.
V-B3 Sensitivity to road metrics accuracy
@@
We envision that road traffic measurements can directly be fed into the learning algorithms as smart cities and roads are equipped with detectors that can provide such critical information. Another alternative is to use road traffic predictions instead of measurements. The machine learning research community has seen a surge of deep learning methods that tackle this prediction task with several solutions that achieve high levels of forecasting accuracy.333 An advantage of such an approach is that the mobile cellular prediction algorithms can receive early information about the vehicular flow and speed that will be evidenced in the BS during the next time slot. This in advance received data can increase further the prediction accuracy of the algorithms. In fact, our preliminary results validate this concept under similar setting as those we reported above and experimented with.
@@
To assess the sensitivity of the mobile cellular predictions to estimation errors in the road predictions, we introduce an estimation error in the flow measurements by the PeMS detectors. In particular, we assume a prediction error of $5\%$ in the number of vehicles per 5-minute intervals. The speed, although time correlated with the flow, especially during rush hours, is not as strongly correlated with the mobile cellular traffic for the majority of the BS. Hence, we do not model speed prediction errors. We add Gaussian noise to the real measurements obtained via PeMS. In particular,
@@
where $v$ denotes the real PeMS flow data and $\hat{v}$ is the estimated flow. Whereas our learning model is fed with the estimated value of the flow variable, we generated the mobile cellular traffic with the true PeMS measurement.
@@
The results are reported in Table IV and show that although the noise in the flow variable decreases cell load prediction accuracy, employing road data is largely beneficial yet. Overall, across all error measures and all BSs and the 4 data sets, the prediction error improvement is between $5.2\%$ and $49.9\%$ for the median when using estimated (with errors) road traffic time series data.
@@
Employing flow, speed and calls—when flow is reported with errors—leads to a reduction in the prediction errors even when contrasted with using new, handover and calls data set: the improvement is between $6.4\%$ and $29\%$ for the medians among the 4 error measures we use.
VI Conclusion
In the past, mobile cellular data have been used to model population dynamics and to enrich our understanding of urban mobility. We take a different look and use population dynamics in a highway context to forecast mobile cellular traffic load. By employing vehicular flow and speed time-series observed in the base stations serving a highway, the traffic load prediction error at each base station can be largely and consistently reduced: between about $5\%$ and $56.5\%$ depending on the error measure, base station and specific data sets. This is a powerful result that can be employed in deep learning models such as graph neural networks for mid- to long-term cellular load forecasting or federated learning for short-term predictions.

### v2

@@ V-A3 Machine learning model
Our model has a single LSTM layer with 16 cells. In the training phase, we use the root mean square propagation and mean squared error (MSE) loss for optimizing the models parameters. We set the length of the historical traffic series to six samples (30 min) due to the speed with which changes in the traffic volume occur in 5-minute intervals. The prediction horizon is the next 5-minute interval. The 24 weeks of data are split 12:6:6 chronologically into training, validation, and testing after which data are shuffled.
V-B Evaluation methodology and protocol
Since our goal is to assess the effectiveness of employing data intrinsic to mobile cellular volume generation, we contrast the prediction performance of the implemented machine learning model when cellular KPIs are used (the baseline) with the model’s performance when these same data are enriched with vehicular flow and speed metrics. We evaluate the predictions over a comprehensive set of highway and cellular conditions. Similar methodology of assessing the impact of data on prediction accuracy is followed in [19] when examining the efficacy of incorporating an urban knowledge graph.
We measure the forecasting performance with the mean absolute error (MAE), mean absolute percentage error (MAPE), MSE and root mean squared error (RMSE). To have a common basis for comparison and to fairly attribute the performance deviation to the data set employed in learning, we train and evaluate the model with the same hyper-parameters Section V-A3 using data from the same period and location but with different features: either containing purely network metrics or a set of network and road traffic metrics. We measure the improvement in prediction as the percentage difference in error by ($ErrM_{net}-ErrM_{net\&road})/ErrM_{net}$, where $ErrM$ is the error measure.
TABLE III: Prediction Performance on Calls Dataset and Flow, Speed and Calls Dataset
@@ V-B Evaluation methodology and protocol
TABLE IV: Prediction Performance on New, Handover, and Total Calls Dataset and the Same Dataset but Enriched with Flow and Speed Time-Series
@@
TABLE V: Prediction Performance with Gaussian Noise Added to the Flow
@@
Refer to caption Refer to caption
Figure 3: Improvement in prediction when employing (left) calls vs calls, flow and speed data and (right) new, handover and total number of calls vs the same data set enriched with flow and speed statistics.
Refer to caption Refer to caption Refer to caption
Figure 4: Prediction error reduction when contrasting: (left) calls (baseline) vs the same data (C) augmented with new, handover and calls (NHC); (center) flow, speed, calls (FSC) (baseline) vs flow, speed, new, handover and total number of calls (FSNHC). (right) NHC data set (baseline) vs FSC.
Refer to caption Refer to caption Refer to caption
Figure 5: Improvement in prediction when the flow is estimated with errors. Contrasting: (left) calls (baseline) vs the same data (C) augmented with vehicular flow estimated with errors and speed ($\hat{F}$); (center) NHC (baseline) vs $\hat{F}$. (right) NHC (baseline) vs $\hat{F}$.
V-C How much can population dynamics data help?
V-C1 Overall performance
@@
Comprehensive results from the four different data sets—1) calls (denoting total number of calls, that is, the sum of new and handover calls) only; 2) flow, speed and calls; 3) new, handover, calls and 4) flow, speed, new, handover, calls—when the call duration is modeled by a mixture of two log-normal distributions are summarized in Table III and Table IV. For each error metric we show the minimum, median and maximum of the 5 simulation runs conducted for each data set and BS. We make the observations as follows:
@@
  • Employing data that captures the processes intrinsic to mobile cellular traffic generation consistently improves cellular load prediction performance. All BSs observe improvements in all their forecasting measures when learning from both network and road traffic data.
@@
    The prediction error reduction among all BSs is between $8.4\%$ and $33.9\%$ (MAE), $13.3\%$ and $56.5\%$ (MSE), $12.5\%$ and $38.4\%$ (MAPE), and $6.9\%$ and $34\%$ (RMSE) when in addition to (total number of) calls, also flow and speed are employed, and when comparing the medians of all BSs, see Figure 3 (left). Similar improvement trends are recorded for the minimum and maximum errors, Table I. Using more refined network KPIs—those that contain the number of new and handover calls, and their total—and road data, reduces the prediction errors too: between $7.5\%$ and $21.8\%$ (MAE), $11\%$ and $35.4\%$ (MSE), $10.9\%$ and $25.7\%$ (MAPE), and $5.7\%$ and $19.7\%$ (RMSE) for the medians, see Figure 3 (right), and likewise for the smallest and largest errors.
@@
  • Road traffic metrics can capture the processes underlying mobile cellular load generation. We employ flow and speed as a means to characterize the underlying population dynamics in highways and through them the data generation process. The flow captures the number of potential call generation sources. The speed can indicate density dynamics and can also serve as a gauge for road traffic congestion and hence, for increased load on the BS. The number of handovers (or equivalently, the handover rate) is an indicator of mobility too yet our results are not conclusive about its impact. Contrasting the performance of calls to that of new, handover, and calls, we observe an error improvement in some of the error measures: it is in the order of no more than $3\%$ for the majority of the BSs, see Figure 4 (left and center plots).
@@
    Finally, employing flow, speed and calls (3 variables) leads to a large reduction in the prediction errors even when contrasted with employing new, handover and calls (3 variables): $8.6\%$ and $22.8\%$ (MAE), $14.2\%$ and $36.3\%$ (MSE), $11.5\%$ and $25\%$ (MAPE), and $7\%$ and $20.2\%$ (RMSE) for the medians, Figure 4 (right plot).
@@
  • Choice of data (split) and hyper-parameter tuning. The shown results are from using 24 weeks of data and 75:25 ratio of train and validation to test data. We observed cases when the performance of some of the BSs is decreased when using road data. This is the case for the MSE and RMSE of BS 320287 when the data set consists of weeks 15 to 21 and the data is chronologically split  into 3:2:2 for training, validation and testing [^1] See https://github.com/nvassileva/DataMatters for detailed results.. We identify week 15 as outlying from the other training weeks and consequently, as the main explanation for the results. For the same period but with 4:1:2 partition of the data, the results are consistent with what we report above—decreased prediction errors across all performance measures and BSs when employing flow and speed time-series in addition to call time-series data. This 4:1:2 data split also turns into smaller forecasting errors for the purely network data sets. In other words, choice of data that captures the underlying patterns and an appropriate hyper-parameter tuning are always required. In our on-going work we are studying the effect of the data split on the overall forecasting performance.
V-C2 Sensitivity to road metrics accuracy
@@
We envision that road traffic measurements can directly be fed into the learning algorithms as smart cities and roads are equipped with detectors that can provide such critical information. Another alternative is to use road traffic predictions instead of measurements. The machine learning research community has seen a surge of deep learning methods that tackle this prediction task with several solutions that achieve high levels of forecasting accuracy [20].222 An advantage of such an approach is that the mobile cellular prediction algorithms can receive in advance information about the vehicular flow and speed that will be evidenced in the BS during the next time slot. This timely received data can increase further the prediction accuracy of the algorithms. In fact, our preliminary results validate this concept under similar conditions as those we report above.
@@
To assess the sensitivity of the mobile cellular predictions to estimation errors in the road predictions, we introduce an estimation error in the flow measurements by the PeMS detectors. In particular, we assume a prediction error of $5\%$ in the number of vehicles per 5-minute intervals, modeled by adding Gaussian noise to the real PeMS measurements:
@@
where $v$ denotes the real PeMS flow data and $\hat{v}$ is the estimated flow. Whereas our learning model is fed with the estimated value of the flow variable $\hat{v}$, we generated the mobile cellular traffic load with the true PeMS measurement $v$.
@@
The results are reported in Table V and show that although the noise in the flow variable decreases cell load prediction accuracy, employing road data is largely beneficial yet. Overall, across all error measures and all BSs and the 4 data sets, the prediction error improvement is between $5.2\%$ and $49.9\%$, Figure 5 (left and center plots), for the medians when using estimated (namely, with errors) road traffic time-series data.
@@
Employing flow, speed and calls—when flow is reported with errors—leads to a reduction in the prediction errors even when contrasted with new, handover and calls data set: the improvement is between $6.4\%$ and $29\%$ for the medians among the MAE, MAPE, MSE, and RMSE error measures, Figure 5 (right plot).
VI Novelty and impact
Novelty. Mobile cellular load forecasting is native to network resource optimization and delivery of services with reliability, latency and quality guarantees. Therefore, it has remained prominent across all generations of communication networks. In the latest research in this area, the primary focus is on developing larger and more powerful machine learning models [1], [2]. When it comes to data, the state-of-the-art contributions (see [2] for an overview) focus on incorporating exogenous information. In particular, these studies rely on environmental data—in addition to cellular time-series—as such auxiliary information describes the BS context and can bring improvements in terms of reduced prediction errors [3], [4]. A few deep learning algorithms [21], [22] explicitly incorporate handover rates between BSs in an attempt to better reflect traffic fluctuations via user mobility. Nonetheless, none of the prior art proposals explores data that can implicitly or explicitly characterize the data sources and their variability. The novelty of our approach consists of modeling the internals of the cellular traffic generation phenomenon. We do this by counting the potential sources of cellular traffic and by accounting for their velocity.
Impact. Employing population dynamics has the potential not only to reduce the uncertainty about future cellular volumes but also to reduce the cost of collecting big and diverse data and hence, to reduce the computational complexity and memory requirements and ultimately, to decrease the energy expenditure too. Lighter learning models are also more versatile as they can be used at the edge and in roadside units too. Furthermore, concept drift [1] arising from changes in network configurations or conditions and users’ behavior would have a major impact on models relying on knowledge graphs as these graphs would need to be updated first.
Potential limitations. The concept of employing population dynamics together with relevant mobile cellular data to accurately predict cellular traffic is of a much broader scope than a specific scenario. We validate our idea in a highway scenario yet we conjecture that it is applicable to the urban setting too. At present, a limitation for the implementation of our idea in practice is the potential lack of real-time population density and mobility data. Currently, real-time road traffic measurements are an integral part of the urban transportation management systems and these can be used in forecasting the service load placed by vehicles and their passengers on the cellular network. In the future, different means for estimating pedestrian density will be available too. The application of optical cables for simultaneous data transmission and tracking of people in a train station, for instance, is discussed in [23].
VII Conclusion
Mobile cellular data have been used to model population dynamics with the aim to better understand urban mobility. We take a different look and employ population dynamics in a highway to forecast mobile cellular traffic on the highway. Our contributions stem from proposing the type of data that can, in practice, substantially scale up the accuracy of mobile cellular traffic predictions. In particular, we propose to employ data that characterizes the sources of communication volume and its dynamics in addition to BS traffic load.
We validate our hypothesis in a real highway scenario by using vehicular flow and speed time-series. The results show that the traffic load prediction error at each BS can be largely and consistently reduced: between about $5\%$ and $56.5\%$ depending on the error measure, BS and specific data sets. These results are obtained for BSs featuring realistic conditions including different range, fluctuating road traffic and cellular volumes.
We sincerely hope that this powerful result will prompt future research. Our proposal can boost the prediction accuracy of centralized and distributed learning models for short-, mid- and long-term cellular load forecasting. The idea of using population dynamics together with relevant mobile cellular data, is of much broader scope than a specific scenario. Examining and quantifying the reductions in prediction error with mobile operator and road data on highways and urban scenarios would be the next step.